......................................................................
A man has 10 gallons of a 50% sulphuric acid mixture, and 20 gallons of a 20% solution, and 5 gallons of a 10% solution. He wants to use up all the 10% solution and make 15 gallons of 30% solution. How much of each solution should he use?

Answer by stanbon(75887)

A man has 10 gallons of a 50% sulphuric acid mixture, and 20 gallons of a 20% solution, and 5 gallons of a 10% solution. He wants to use up all the 10% solution and make 15 gallons of 30% solution. How much of each solution should he use?
---------------
Equation:
acid + acid + acid = acid
-----
0.10*5 + 0.50*x + 0.20(10-x) = 0.30*15
-------
Multiply thru by 100 to get:
10*5 + 50x + 20*10-20x = 30*15
--------
50 + 30x + 200 = 450
30x = 200
x = 20/3 = 6 2/3 gallons (amt. of 50% solution needed)
10-x = 3 1/3 gallons (amt. of 20% solution needed)
==============
Cheers,
Stan H.
==============